Samantha’s student loans total $20,000. Part was a personal loan at 12% interest; the other was a Stafford loan at 10%. After one year the loans accumulated $2160 in interest. What was the amount of each loan?
Answer by stanbon(57307)

Samantha’s student loans total $20,000. Part was a personal loan at 12% interest; the other was a Stafford loan at 10%. After one year the loans accumulated $2160 in interest. What was the amount of each loan?
----------------------------
Equations:
Quantity: x + y = 20,000
Interest:0.12x+0.10y = 2150
-----------------------------------
Multiply thru the Quantity equation by 12
Multiply thru the Interest equation by 100
-----------------------------------
12x + 12y = 12*20000
12x + 10y = 215000
--------------------------
Subtract the 2nd equation from the 1st and solv for "y":
2y = 25000
y = $12,500 (amt. invested at 10%)
x = 20,000-12,500 = $7,500 (amt. invested at 12%)
====================================================
Cheers,
Stan H.